### v1

# Integer Lattice Gas with a sampling collision operator for the fluctuating Navier-Stokes Equation

Abstract
While lattice Boltzmann methods were originally derived as the Boltzmann limit of a Boolean lattice gas, modern improvements to lattice Boltzmann methods have broken this link. Because of this most lattice Boltzmann methods are not unconditionally stable and extending them to model fluctuations is challenging. In this article we show that the integer lattice gas method developed by Blommel and Wagner has a lattice Boltzmann limit corresponding to entropic lattice Boltzmann methods, while also recovering proper ideal gas equilibrium fluctuations. We also show that it is possible to accelerate these integer lattice gases such that they become computationally competitive with the corresponding lattice Boltzmann method.
Lattice Gas, Lattice Boltzmann, Monte Carlo, Fluctuatons
I Introduction
Frisch, Hasslacher and Pomeau’s Frisch et al. (1987) extension of lattice gas automata to allow for the simulation of hydrodynamics opened up the new field of discrete simulations of fluid dynamics. Despite their very simple local evolution rules these lattice gases could be shown to recover the continuity and Navier-Stokes equations. The fluctuations of these lattice gases closely resembled the fluctuations of real liquids Rivet and Boon (2001).
Initially the Boltzmann limit was applied to lattice gas systems to analyze their hydrodynamic limit in analogy to the procedure in kinetic theory Huang (1987). This Boltzmann limit consists of taking an ensemble average that removes the fluctuations. McNamara et al. McNamara and Zanetti (1988) were the first to use this Boltzmann limit of the lattice gas as a simulation method in its own right. Because simulating the lattice Boltzmann equation eliminated statistical noise, this was an advantage for simulating noise-free hydrodynamics. Previously this had to be achieved by performing a statistical average over several lattice gas simulations.
The lattice Boltzmann method was later decoupled from the lattice gas entirely when Qian, d’Humieres, and Lallemand Qian et al. (1992) proposed a simplified BGK collision operator. By simply choosing a local equilibrium distribution (the now ubiquitous second order polynomial equilibrium distribution) they simplified the algorithm and made it significantly more accessible. They were also able to remove some non-Galilean invariant artifacts of Boolean lattice gas systems, such as a velocity dependent viscosity Frisch et al. (1987); Higuera and Jimenez (1989) and pressure, even if these deviations became most notable in parameter ranges (i.e. at high Mach numbers) that were still not accessible to these lattice Boltzmann methods.
The simplicity of this BGK collision operator, more than anything else, influenced the explosive rise of lattice Boltzmann methods over their ancestral lattice gas methods. Certain systems, however, such as the Brownian motion of colloids considered by Ladd et al. Ladd et al. (1988), require fluctuations in order to be simulated. These fluctuating systems were, at the time, limited to being simulated by lattice gases. To allow these simulations to benefit from the advances of lattice Boltzmann methods, Ladd developed a fluctuating lattice Boltzmann method. The basic idea was to implement Landau and Lifshitz’ Landau et al. (1959) fluctuating stress tensor into the collision operator leading to a Langevin approach Lemons and Gythiel (1997) for the fluctuating lattice Boltzmann method. This fluctuating lattice Boltzmann method recovered the fluctuating hydrodynamic equations Ladd (1993). However, utilizing the fluctuating stress tensor alone limited the recovery of the correct fluctuations to those with a wavelength on the order of the system size, while the total order of fluctuations was suppressed by about a factor of two.
Adhikari et al. discovered that in order to recover the smaller scale fluctuations, a full set of fluctuating modes, rather than just those associated with hydrodynamic quantities, was needed Adhikari et al. (2005). This method is restricted to ideal gas systems. For non-ideal systems additional corrections are required Belardinelli et al. (2015). Since the Langevin approach is based on continuous densities, the large fluctuations at low densities will lead to the formation of non-physical negative densities (Kähler, 2012, sec. 5.3).
When considering methods for simulating fluctuating systems, however, it is much more natural to think in terms of discrete systems since fluctuations fundamentally arise because of the discreteness of nature Einstein (1906). This inspired Chopard et al. to instead develop novel lattice gas algorithms instead of adding fluctuations artificially to the lattice Boltzmann algorithm (Chopard et al., 1998; Chopard and Droz, 1998, sec. 5.7). They extended the occupation numbers of the lattice gas from Boolean values to integer values, a concept that was first introduced by Chen et al. the year prior with their Digital Physics approach Chen et al. (1997). This allowed for the lattice gas system of Chopard et al. to recover a Poisson distribution in global equilibrium. Secondly, Chopard et al. implemented a sampling algorithm, which was originally theorized by Boghosian et al. Boghosian et al. (1997), that distributed particles randomly across the system velocities based on a local equilibrium ensemble. For the diffusive case, a multinomial distribution over the occupation numbers themselves was utilized for the local equilibrium ensemble, which was sufficient to follow the conservation of mass required by the system (Chopard and Droz, 1998, sec. 5.7).
In extending the method to hydrodynamic systems, however, Chopard et al. did not utilize an explicit local equilibrium ensemble due to the complexity involved in deriving this distribution while accounting for momentum conservation. Instead, a multinomial distribution similar to that of the diffusive case was used, centered around the second order polynomial equilibrium distribution. This on its own would violate momentum conservation, so a random factor that redistributed a single particle at a time until momentum conservation was upheld was also added Chopard et al. (1998). Both of these methods were, for higher densities, well approximated by Gaussian distributions, which caused their algorithmic complexity to be approximately O(1) with respect to the system density, matching the complexity of the lattice Boltzmann algorithm.
Blommel and Wagner created another integer based lattice gas algorithm to simulate hydrodynamic systems Blommel and Wagner (2018). Their algorithm utilized a series of random two-particle collisions in order to redistribute particle velocities while still maintaining momentum conservation with collision rules explicitly derived by imposing detailed balance on the system. This allowed Blommel and Wagner to recover Poisson distributed fluctuations in global equilibrium similar to Chopard et al..
When analyzing their algorithm, the global equilibrium average distribution was found to correspond not to the polynomial lattice Boltzmann equilibrium distribution, but to the entropic lattice Boltzmann equilibrium distribution. This equilibrium distribution had already been derived by Ansumali et al. Ansumali et al. (2003) previously by minimizing an H-functional. The polynomial equilibrium distribution utilized by Chopard et al. Chopard et al. (1998) for their hydrodynamic multiparticle lattice gas, therefore, was actually incompatible with a microscopic lattice gas implementation.
Since the number of individual two-particle collisions needed to achieve a state close to local equilibrium scales with the square of the number of particles at each lattice site, Blommel and Wagner’s algorithm proved much slower than the fluctuating lattice Boltzmann algorithm even for moderate local densities of a few hundred particles per lattice site. This lead Seekins and Wagner Seekins and Wagner (2022) to accelerate Blommel’s algorithm. A sampling method similar to that of Chopard et al. was implemented in order to reduce the number of random numbers needed, leading to increased efficiency.
Using detailed balance in order to derive a local equilibrium distribution, Seekins and Wagner recovered an identical multinomial distribution to that of Chopard et al. (Chopard and Droz, 1998, sec. 5.7). Although this meant that the diffusive algorithm created by Seekins and Wagner was essentially identical to that of Chopard et al., the derivation of the equilibrium distribution via detailed balance was novel, and could be used to derive the proper equilibrium distribution for hydrodynamic systems. The efficiency of the algorithm was comparable to a diffusive fluctuating lattice Boltzmann algorithm, along with having a complexity at high densities of approximately O(1) with respect to density Seekins and Wagner (2022); Wagner and Strand (2016).
In this paper, we extend our prior work for diffusive systems to hydrodynamic systems by deriving an appropriate local equilibrium distribution. for simplicity, we limit the system to one dimension, leaving an extension to higher dimensions to future work. The paper is structured as follows: in Section II of the paper, we will cover the sampling lattice gas algorithm in detail. In Section III we derive the local equilibrium distribution and provide an analysis of the Poisson distribution as a global equilibrium distribution for the integer lattice gas. Section IV will discuss the Boltzmann limit of the lattice gas and its uniqueness. Section V covers the results of two test cases used to study the non-equilibrium dynamics of the system. Finally, section VI will compare the efficiency of our sampling lattice gas to that of both Blommel and Wagner and an entropic lattice Boltzmann algorithm.
II The Sampling Lattice Gas Algorithm
II.1 Sampling Lattice Gas Basics
A lattice gas consists of a set of discrete cells situated on a lattice, each containing a discrete number of velocity states $v_{i}$ for particles to occupy Wolf-Gladrow (2004). The velocities are usually identical for every cell, and are defined such that if $x$ is a position on the lattice, then for any velocity $v_{i}$, $x+v_{i}$ is also a position on the lattice, assuming a timestep of $\Delta t=1$. These velocities are also typically restricted in number such that the particle displacement over a single timestep is small. The particles are accounted for by the occupation numbers $n_{i}(x,t)$, denoting the particles that move from lattice site $x-v_{i}$ at time $t-1$ to lattice site $x$ at time $t$. These particles are then redistributed among the occupation numbers, observing the applicable conservation rules, and streamed. This two-step process represents a discrete timestep for the system. Thus, the $n_{i}(x,t)$ evolve according to,
$n_{i}(x+v_{i},t+1)=n_{i}(x,t)+\Xi_{i}[\{n_{j}\}_{g}],$ (1)
which is known as the lattice gas evolution equation. Here $\{n_{j}\}_{g}$ represents the set of all $n_{i}$ of the system at the current time. In this evolution equation, the collision operator $\Xi_{i}$ is a stochastic operator. The explicit nature of the $n_{i}$ is dependent on the lattice gas system observed. For Boolean lattice gases, for example, these $n_{i}$ would be Boolean in nature, leading to $\Xi_{i}$ to be calculable explicitly within a lookup table due to the limited nature of the probability space for single-occupation velocity states Frisch et al. (1986, 1987). In our integer lattice gas method, the $n_{i}$-values become integer based, which leads to a collision operator $\Xi_{i}$ that can take on integer values.
In the ideal gas systems considered here the collision operator will only act on the local distribution of occupation numbers, $\{n_{j}\}_{l}$, which consists only of the local occupation numbers at position $x$ and $t$. In this nomenclature we exclude the implicit dependence on $x$ and $t$ to make the equations more readable. However, in principle the collision operator can depend on the global distribution $\{n_{j}\}_{g}$, as would be the case for non-ideal lattice gas systems. An example of that would be Rothmann and Keller’s algorithm Rothman and Zaleski (1994) that is now known as the color gradient method. This method depends on neighboring site average density gradients.
The sampling collision operator takes inspiration from the BGK collision operator for the Boltzmann equation. The Boltzmann equation is an evolution equation for a distribution function $f_{i}(x,t)$, and the BGK collision operator is an approximation that simply moves the system towards it’s local equilibrium distribution $f_{i}^{0}(x,t)$ Bhatnagar et al. (1954) where the local equilibrium distribution depends on the locally conserved quantities. This BGK lattice Boltzmann algorithm will be discussed in more detail in upcoming sections.
In a lattice gas system, the system is described by $\{n_{i}\}_{g}$. Unlike in the lattice Boltzmann case, the local equilibrium state for the $n_{i}$ is not given by an equilibrium distribution, but rather by a probability distribution around some average value, known as the local equilibrium ensemble. This local equilibrium ensemble, given a local pre-collision set of occupation numbers $\{n_{i}\}_{l}$, gives a probability $P^{0}(\{\hat{n}_{i}\}_{l};\{n_{i}\}_{l})$ for each post-collision set $\{\hat{n}_{i}\}_{l}$. A post-collision set of occupation numbers can be then be sampled from the local equilibrium ensemble in a process discussed in detail in Seekins and Wagner’s paper Seekins and Wagner (2022). The collision operator for the sampling lattice gas algorithm thus becomes the difference between the pre-collision and post-collision occupation numbers, explicitly defined as,
$\Xi_{i}[\{n_{j}\}_{l}]=\hat{n}_{i}-n_{i}.$ (2)
However, this representation of the collision operator does not account for the fact that the system may not relax fully to local equilibrium in one timestep. In order to simulate this under-relaxation in a system of discrete particles, we can represent the fraction of particles that should undergo collision in that timestep as $\omega$. We create a subset of our occupation numbers $\{n^{\omega}_{i}\}_{l}$ to undergo collision, with the probability that any single particle is a member of the subset being given by $\omega$. Therefore, each specific value of $n_{i}^{\omega}$ has a probability associated with it given by the binomial distribution,
$P(n^{\omega}_{i})=\left(\begin{array}[]{c}n_{i}\\ n^{\omega}_{i}\end{array}\right)\omega^{n^{\omega}_{i}}(1-\omega)^{n_{i}-n^{\omega}_{i}}.$ (3)
We are able to obtain the number of particles for each $n_{i}^{\omega}$ by sampling out of this binomial distribution. We create a pre-collision set $\{n^{\omega}_{i}\}_{l}$, and then use the probability distribution $P^{0}(\{\hat{n}^{\omega}_{i}\}_{l};\{n^{\omega}_{i}\}_{l})$ to sample a new, post collision set of occupation numbers $\{\hat{n}^{\omega}_{i}\}_{l}$. The collision operator for a specific value of $\omega$ is then given by,
$\Xi_{i}^{\omega}[\{n_{j}\}_{l}]=\hat{n}^{\omega}_{i}-n_{i}^{\omega}.$ (4)
It remains to derive the local equilibrium ensemble explicitly.
II.2 Derivation of the Local Equilibrium Ensemble
In the hydrodynamic case considered here there are two conserved quantities: mass and momentum. Since momentum is conserved, the probability distribution for the $n_{i}$ will be correlated. However, by transforming our representation of the system from being expressed as an set of individual occupation numbers $\{n_{i}\}_{l}$ into being expressed by the hydrodynamic moments we may be able to obtain an uncorrelated probability distribution. These moments are physical quantities within the simulation, with a total number given by the total number of system velocities $Q$. From some fundamental particle collision process like that of Blommel and Wagner Blommel and Wagner (2018) we deduce the local equilibrium probability distributions for the non-conserved moments. We then use these local equilibrium ensembles to sample the post-collision moments as random variables. This is the extent of how far we are able to describe our algorithm in all generality.
In order to know what moment representation is applicable to our system, we must focus our attention on a specific example. As this paper serves as a proof of concept for our sampling algorithm, we focus on the simplest possible system that recovers hydrodynamic effects, a one-dimensional system with three total velocity vectors, leaving the extension of the algorithm to more complex systems, including higher dimensional systems, to future work.
Also known as a D1Q3 system, the velocity vectors of this system connect a given lattice point to its nearest neighbors, and include a zero velocity state. Numerically, we can represent this by restricting our velocities to the set $v_{i}\in\{-1,0,1\}$, with indices $i$ that are given by the value of the velocity they represent (i.e. $v_{-1}=-1$). Knowing the structure of our system, we know that we have a total of three moments due to having a total of three velocities, and because two of our local system quantities are conserved, their corresponding moments must be conserved as well. Thus, we can define the local density $N$ and the local momentum $J$ as moments,
$N$ $=n_{1}+n_{0}+n_{-1},$ (5)
$J$ $=n_{1}-n_{-1}.$ (6)
We define a third moment $\pi$ as,
$\pi=n_{1}+n_{-1},$ (8)
which, not being a conserved quantity, is the only moment that is allowed to change over the course of a local collision step. Each combination of these three moments fully describes a unique occupation state $\{n_{i}\}_{l}$, as Eqs. (5), (6), and (8) can be inverted to recover our system’s occupation numbers,
$n_{1}$ $=\frac{\pi+J}{2},$ (9)
$n_{0}$ $=N-\pi,$ (10)
$n_{-1}$ $=\frac{\pi-J}{2}.$ (11)
Thus, we can redefine $P^{0}(\{\hat{n}_{i}\}_{l};\{n_{i}\}_{l})$ in terms of $N$, $J$, and $\pi$ as $P^{0}(\hat{N},\hat{J},\hat{\pi};N,J,\pi)$, or the probability of finding a system in a state with moments $\hat{N}$, $\hat{J}$, and $\hat{\pi}$ after a collision given a pre-collision state with moments $N$, $J$, and $\pi$. We can then simplify this representation, as our conservation rules state that $\hat{N}=N$ and $\hat{J}=J$. Furthermore, because we are assuming that the system is in local equilibrium when described by this distribution, we know that $\hat{\pi}$ does not depend on the pre-collision $\pi$-value. The probability distribution for our ensemble can then be written as $P^{0}(\pi;N,J)$, which is the equilibrium probability of finding the system in a post-collision state with a $\pi$-value of $\pi$, given a pre-collision system with conserved moments $N$ and $J$, and an arbitrary initial value for $\pi$.
We now derive the particular probability distribution that recovers the integer lattice gas results obtained by Blommel and Wagner. In local equilibrium, this distribution must be invariant under collisions. By using Blommel and Wagner’s algorithm as a basis, we restrict the possible collisions of our system to only involving pairs of particles. These collisions conserve momentum, and their equilibrium ensemble must be stationary over each collision. Thus, we can look at the equilibrium distribution that corresponds to Blommel and Wagner’s algorithm as the effect of many individual two-particle collisions.
Due to $P^{0}(\pi;N,J)$ being invariant under a single collision we can write,
$T_{\pi\to\pi+2}P^{0}(\pi;N,J)=T_{\pi+2\to\pi}P^{0}(\pi+2;N,J),$ (12)
where $T_{\pi\to\pi+2}$ and $T_{\pi+2\to\pi}$ are the transition probabilities between states $\pi$ and $\pi+2$ under a single two-particle collision. These transition probabilities represent the probability of any collision occurring that increases or decreases the local $\pi$-value respectively. In order to write these transition probabilities explicitly, we take the sum of the probabilities of each possible collision that increases or decreases the $\pi$-value depending on which transition probability we wish to write. There are only four two-particle collisions for a D1Q3 system that alter the $\pi$-value, two that increase the $\pi$-value by two and two that decrease it by two,
$\pi\rightarrow\pi+2\biggr{\{}\begin{array}[]{c}\{0,0\}\to\{1,-1\}\\ \{0,0\}\to\{-1,1\},\\ \end{array}\hskip 5.69046pt$ (13)
$\pi+2\rightarrow\pi\biggr{\{}\begin{array}[]{c}\{-1,1\}\to\{0,0\}\\ \{1,-1\}\to\{0,0\}.\\ \end{array}\hskip 8.5359pt$ (14)
The probabilities of each of these possible collisions occurring is given by the probability of picking a pair of particles with the proper set of velocity indices $\{i,j\}$, which is written as $p_{\{i,j\}}$, multiplied by the probability of the collision occurring from velocity state $\{i,j\}$ to velocity state $\{k,l\}$, written as $P_{\{i,j\}\to\{k,l\}}$. Thus, $T_{\pi\to\pi+2}$ and $T_{\pi+2\to\pi}$ can be defined as,
$T_{\pi\to\pi+2}$ $=p_{\{0,0\}}P_{\{0,0\}\to\{1,-1\}}+p_{\{0,0\}}P_{\{0,0\}\to\{-1,1\}},$ (15)
$T_{\pi+2\to\pi}$ $=p_{\{1,-1\}}P_{\{1,-1\}\to\{0,0\}}+p_{\{-1,1\}}P_{\{-1,1\}\to\{0,0\}}.$ (16)
We assume that each particle has an equal likelihood to undergo collision. Thus, $p_{\{i,j\}}$ can be defined by considering the fraction of particles in the relevant $n_{i}$ occupation states compared to the total density at their lattice site $N$,
$p_{\{i,j\}}=\frac{n_{i}(n_{j}-\delta_{ij})}{N(N-1)},$ (17)
where $\delta_{ij}$ is the Kronecker delta, and is incorporated to avoid double counting particles should the indices $i$ and $j$ be the same. The $P_{\{i,j\}\to\{k,l\}}$-values were defined explicitly by Blommel and Wagner such that at zero velocity the typical lattice Boltzmann equilibrium values are recovered on average Blommel and Wagner (2018). For the relevant collisions in Eqs. (15) and (16), these values are given by,
$P_{\{0,0\}\to\{1,-1\}}=1/16,$ (18)
$P_{\{0,0\}\to\{-1,1\}}=1/16,$ (19)
$P_{\{1,-1\}\to\{0,0\}}=1,$ (20)
$P_{\{-1,1\}\to\{0,0\}}=1.$ (21)
It is not obvious that the probability $P_{\{i,j\}\to\{k,l\}}$ should necessarily be independent of the locally conserved quantities. However, as Blommel and Wagner’s method makes this assumption, we make it as well in order to adapt their algorithm. A true test of the validity of this approach could be done by looking at the coarse-graining of a molecular dynamics simulation, similar to the work done by Parsa et al. Parsa and Wagner (2017).
We can now rewrite the transition probabilities as,
$T_{\pi\to\pi+2}$ $=\frac{n_{0}(n_{0}-1)}{8N(N-1)},$ (22)
$T_{\pi+2\to\pi}$ $=\frac{2n_{1}n_{-1}}{N(N-1)},$ (23)
and transform them into moment representation by using Eqs. (9), (10), and (11),
$T_{\pi\to\pi+2}$ $=\frac{(N-\pi)(N-\pi-1)}{8N(N-1)},$ (24)
$T_{\pi+2\to\pi}$ $=\frac{\left((\pi+2)^{2}-J^{2}\right)}{2N(N-1)}.$ (25)
Eq. (12) can then be rewritten as,
$\frac{(N-\pi)(N-\pi-1)}{8N(N-1)}P^{0}(\pi;N,J)=$ (26)
$\frac{\left((\pi+2)^{2}-J^{2}\right)}{2N(N-1)}P^{0}(\pi+2;N,J).$
Simplifying this expression gives us a recursive formula for $P^{0}(\pi;N,J)$,
$P^{0}(\pi+2;N,J)=\frac{(N-\pi)(N-\pi-1)}{4((\pi+2)^{2}+J^{2})}P^{0}(\pi;N,J).$ (27)
This recursive formula, along with the requirement for normalization,
$\sum_{\pi}P^{0}(\pi;N,J)=1,$ (28)
fully describes the local equilibrium ensemble. If we sample random $\hat{\pi}$-values out of this local equilibrium ensemble, our collision operator in moment space becomes,
$\Xi_{\pi}=\hat{\pi}-\pi.$ (29)
This representation of the collision operator is entirely equivalent to the representation in Eq. (2). Due to the required momentum conservation the local equilibrium ensemble will, starting at $\pi=J$, increment by two, and thus will maintain its initial parity. Initial $N$- and $J$-values of opposing parity result in a maximum possible $\pi$-value of $N-1$. Also, since the particles occupying an $n_{0}$ state do not stream between lattice sites, and are only altered in multiples of two, the parity of the $n_{0}$-value at each lattice site from initialization acts as a spurious conserved quantity, the (minor) consequences of which will be discussed in the upcoming sections.
The explicit formula corresponding to the recursive formula of our local equilibrium ensemble can be found by considering the evolution of the correct global equilibrium distribution, something that will be discussed at length in the next section, including a derivation of this explicit formula. This explicit formula is given by,
$\frac{1}{P^{0}(\pi;N,J)}=4^{N-\pi}\left(N-\pi\right)!\left(\frac{\pi+\left|J\right|}{2}\right)!\left(\frac{\pi-\left|J\right|}{2}\right)!$ (30)
$\sum_{k=0}^{\left\lfloor\frac{N-\left|J\right|}{2}\right\rfloor}\left(4^{N-2k-\left|J\right|}\left(N-2k-\left|J\right|\right)!\left(k+\left|J\right|\right)!\left(k\right)!\right)^{-1},$
where $k$ is an index which enumerates the possible values of $\pi$ while taking into account the fact that those values increment by multiples of two, and the floor function $\lfloor$..$\rfloor$ … ⌋ returns the greatest integer value less than its argument.
Figure 1: The Local equilbrium ensemble of $\pi$-values for $N=9$ and $J=0$. A single lattice site was made to undergo 100,000 collisions for both Blommel and Wagner’s algorithm and our sampling algorithm  Blommel and Wagner (2018). Then, a further 100,000,000 collisions were run, collecting the $\pi$-value after each. The histograms of these data sets are compared to the recursive and explicit formulae given by Eqs. (27) and (30) respectively
A comparison between this explicit formula and its recursive counterpart can be seen in Fig. 1, along with histograms of local equilibrium $\pi$-values from both Blommel’s algorithm and the sampling algorithm put forward in this paper. A low density system with a local $N$-value of $9$ was chosen due both to the fact that this local equilibrium ensemble is more distinct from a binomial distribution at lower densities and to the fact that we wished to display the lack of negative densities present in the distribution.
The histograms of data from both Blommel and Wagner’s algorithm and our sampling lattice gas algorithm show excellent agreement to the recursive and explicit formulae of the theoretical local equilibrium ensemble shown in Eqs. (27) and (30), as well as maintaining the proper parity for their initial $J$-value. To ensure proper agreement, additional tests were run for different configurations of $N$- and $J$-values, finding results consistent with those shown in Fig. 1. Thus, for brevity, we show only this single example.
II.3 Global Equilibrium Ensemble Analysis and finding an Explicit Formula for the Local Equilibrium Ensemble
Analyzing the global equilibrium ensemble of our lattice gas system serves two primary purposes. Firstly, it allows us to better define our lattice gas algorithm’s equilibrium state for the purpose of discussing the nature of its Boltzmann limit, which we will discuss in more detail in the next section. Secondly, it allows us to analytically derive an explicit formula for the local equilibrium ensemble, defined in the prior section as Eq. (30).
Blommel and Wagner followed Adhikari et al. Adhikari et al. (2005) in using arguments by Landau and Lifshitz (Landau and Lifshitz, 1969, Sec. 114) to postulate that the global equilibrium ensemble of a lattice gas should be given by a Poisson distribution around an equilibrium value Blommel and Wagner (2018). Due to the spurious parity conservation of the $n_{0}$-value at each lattice site this probability distribution cannot be the correct equilibrium ensemble for any specific set of initial conditions. Instead, particular microstates have to be removed from the total Poisson distribution. However, Blommel and Wagner showed numerically that the occupation numbers recovered in equilibrium are approximately Poisson distributed when the initial conditions themselves give a random distribution of even and odd $n_{0}$ and the equilibrium is understood as an ensemble average over these initial conditions.
The Poisson distribution of independent $n_{i}$-values around a given set of means $\tilde{f}_{i}$ is given by:
$P^{\mathrm{Pois}}(\{n_{i}\}_{l},\tilde{f}_{i})=\prod_{i}\exp\left(-\tilde{f}_{i}\right)\frac{(\tilde{f}_{i})^{n_{i}}}{(n_{i})!},$ (31)
where $P^{\mathrm{Pois}}(\{n_{i}\}_{l},\tilde{f}_{i})$ is the probability of finding a particular lattice site in state $\{n_{i}\}_{l}$. The $\tilde{f}_{i}$-values for Blommel and Wagner’s algorithm are given by Blommel and Wagner (2018),
$f_{i}^{0}=$ (32)
$\bar{N}w_{i}\left(1+3v_{i}u+(3v_{i}^{2}-1)\left(\sqrt{1+3(u)^{2}}-1\right)\right),$
which is the equilibrium distribution for the entropic lattice Boltzmann method. The exact formulation of the ensemble average density $\bar{N}$ and velocity $u$ will be discussed in the next section, however in a global equilibrium system they are constant quantities that only depend on the initial conditions of that system,
$\bar{N}^{eq}=\sum_{i}\sum_{x}\frac{n_{i}(x,0)}{L},$ (33)
$\bar{N}^{eq}u^{eq}=\bar{J}^{eq}=\sum_{i}\sum_{x}\frac{n_{i}(x,0)v_{i}}{L},$ (34)
where $L$ is the total lattice length of the system. This means that $f_{i}^{0}(\bar{N}^{eq},u^{eq})$ is a constant quantity.
We can write Eq. (31) in terms of the system moments by utilizing Eqs. (9), (10), and (11). For brevity, the $n_{i}$-values are kept as $n_{i}(N,J,\pi)$ until otherwise required for calculation,
$P^{\mathrm{Pois}}[\{N,J,\pi\}_{l},f_{i}^{eq}(\bar{N}^{eq},\bar{J}^{eq})]=$ (35)
$\prod_{i}\exp[-f_{i}^{eq}(\bar{N}^{eq},\bar{J}^{eq})]\frac{f_{i}^{eq}(\bar{N}^{eq},\bar{J}^{eq})^{n_{i}(N,J,\pi)}}{n_{i}(N,J,\pi)!}.$
Here $u^{eq}$ is replaced by $\bar{J}^{eq}$, as $u^{eq}$ can be written as $\bar{J}^{eq}/\bar{N}^{eq}$.
We can further analyze this Poisson distribution by considering the requirements for equilibrium. Specifically, we require that our global ensemble must be invariant under Eq. (1). As the $n_{i}$-values from Eq. (31) are independently distributed, this distribution is inherently invariant under streaming. Therefore, it only remains to be proven that the system is invariant under collisions. We can test this analytically by considering the effect of a single collision on the probability distribution $P^{\mathrm{Pois}}[\{N,J,\pi\}_{l},f_{i}^{eq}(\bar{N}^{eq},\bar{J}^{eq})]$ as we did in Eq. (12), while adjusting for the fact that the effect of the collision operator is only required to be zero on average. Thus, we can write,
$\sum_{N,J}P^{\mathrm{Pois}}[\{N,J,\pi\}_{l},f_{i}^{eq}(\bar{N}^{eq},\bar{J}^{eq})]T_{\pi\to\pi+2}=$ (36)
$\sum_{N,J}P^{\mathrm{Pois}}[\{N,J,\pi+2\}_{l},f(\bar{N}^{eq},\bar{J}^{eq})T_{\pi+2\to\pi},$
where the sum $\sum_{N,J}$ is a sum over all possible values of $N$ and $J$, with minimum values $N=\pi$ and $J=-\pi$ and maxima of the number of total particles in the system for $N$, and $J=\pi$. Substituting in our previously defined quantities on both sides of this equation gives us,
$\sum_{N,J}\prod_{i}\exp(-f_{i}^{eq})\frac{(f_{i}^{eq})^{n_{i}(N,J,\pi)}}{n_{i}(N,J,\pi)!}\frac{(N-\pi)(N-\pi-1)}{8N(N-1)}=$ (37)
$\sum_{N,J}\prod_{i}\exp(-f_{i}^{eq})\frac{(f_{i}^{eq})^{n_{i}(N,J,\pi+2)}}{n_{i}(N,J,\pi+2)!}\frac{(\pi^{2}-J^{2})}{2N(N-1)}.$
We can further simplify this by taking the product over $i$, and substituting in Eqs. (9), (10), and (11) for the remaining $n_{i}$-values. However, for the sake of clarity, we first define factors $\alpha$ and $\beta$ as common terms on both sides of the equation after the substitution,
$\alpha$ $=\frac{1}{4}\exp(-f_{0}^{eq}-f_{1}^{eq}-f_{-1}^{eq})\frac{1}{N(N-1)},$ (38)
$\beta$ $=\frac{(f_{0}^{eq})^{N-\pi}(f_{1}^{eq})^{\frac{\pi+J}{2}}(f_{-1}^{eq})^{\frac{\pi-J}{2}}}{\left(N-\pi\right)!\left(\frac{\pi+J}{2}\right)!\left(\frac{\pi-J}{2}\right)!}.$ (39)
Thus, Eq. (37) becomes,
$\sum_{N,J}\frac{1}{2}\alpha\beta\left(N-\pi\right)\left(N-\pi-1\right)=$ (40)
$\sum_{N,J}8\alpha\beta\frac{f^{eq}_{1}f^{eq}_{-1}}{(f^{eq}_{0})^{2}}\left(N-\pi\right)\left(N-\pi-1\right).$
This equation can be simplified once more by calculating $f_{1}^{eq}f_{-1}^{eq}/(f_{0}^{eq})^{2}$, which can be calculated using Eq. (32), and comes out to a value of $1/16$. This, when substituted into Eq. (40), makes both sides of the equation identical, which proves that $P^{\mathrm{Pois}}[\{N,J,\pi\}_{l},f_{i}^{eq}(\bar{N}^{eq},\bar{J}^{eq})]$ is invariant under collisions. In fact, because we proved equivalence for each individual term of the sum over $N$ and $J$, the aforementioned Poisson distribution must be a representation of the local equilibrium ensemble for our lattice gas when considered for fixed $N$ and $J$,. Specifically, we find,
$P^{0}(\pi;N,J)=\gamma(\pi)\frac{1}{\sum^{N}_{\pi^{\prime}=|J|}\gamma(\pi^{\prime})}.$ (41)
where $\gamma(\pi)$ is given by
$\gamma(\pi)=\mathrm{exp}\left(f_{0}^{eq}f_{1}^{eq}f_{-1}^{eq}\right)(\frac{f_{0}^{eq})^{N-\pi}(f_{1}^{eq})^{\frac{\pi+J}{2}}(f_{-1}^{eq})^{\frac{\pi-J}{2}}}{(N-\pi)!\left(\frac{\pi+J}{2}\right)!\left(\frac{\pi-J}{2}\right)!}.$ (42)
Because Eq. (42) does not include all possible microstates of the Poisson distribution, it is no longer normalized. To account for this, Eq. (41) includes a normalization factor. This normalization factor consists of a sum over all possible values of $\pi$, utilizing an index $\pi^{\prime}$ to do so. However this representation does not take into account that the $\pi$-value can only be incremented by multiples of two. By defining a second index $k$, and setting $\pi^{\prime}=2k+|J|$, we can find a system that increases by two between values of $\pi^{\prime}$ with a lower bound of 00, and an upper bound of $\lfloor(N-|J|)/2\rfloor$. The floor function takes into account the fact that the maximum value of $\pi$ depends on the relative parity of $N$ and $J$.
Blommel and Wagner make the assumption when defining their collision rules that only local quantities can affect the local collisions Blommel and Wagner (2018). Thus, as a representation of the local equilibrium ensemble of Blommel and Wagner’s algorithm, Eq. (41) must only depend on local quantities. Thus, we must be able to rewrite Eq. (41) such that the dependency on the $f_{i}^{eq}$-values, which depend on the global quantities $\bar{N}^{eq}$ and $\bar{J}^{eq}$, vanishes. By using the aforementioned relation that $f_{1}^{eq}f_{-1}^{eq}/(f_{0}^{eq})^{2}=1/16$ to substitute $f_{0}=\sqrt{16f_{1}^{eq}f_{-1}^{eq}}$, we can rewrite $\gamma(\pi)$ as,
$\gamma(\pi)=$ (43)
$\mathrm{exp}\left[4(f_{1}^{eq}f_{-1}^{eq})^{3/2}\right]\frac{(f_{1}^{eq})^{\frac{N+J}{2}}(f_{-1}^{eq})^{\frac{N-J}{2}}}{4^{N-\pi}(N-\pi)!\left(\frac{\pi+J}{2}\right)!\left(\frac{\pi-J}{2}\right)!}.$
In this form, $\gamma(\pi)$ has all $f_{i}^{eq}$ terms independent of $\pi$, meaning that when the whole of Eq. (41) considered, the $f_{i}^{eq}$ terms will cancel out due to the normalization factor, which, along with making the aforementioned index substitution for $\pi^{\prime}$, recovers Eq. (30).
A collection of these local equilibrium ensembles can then in theory form the proper global equilibrium ensemble for our system, taking into account the spuriously conserved $n_{0}$ parity by picking only $N$- and $J$-values for the global equilibrium ensemble that maintain this conservation.
III The Boltzmann Limit and its Uniqueness
To derive the hydrodynamic limit of our lattice gas algorithm, which will be given by a set of continuity and Navier-Stokes equations, it is required that we first know the Boltzmann limit of our lattice gas. The Boltzmann limit of a lattice gas corresponds to the ensemble average over the possible microstates of a lattice gas system given a specific initial state and elapsed time. The Boltzmann limit of the lattice gas evolution equation then gives us the corresponding lattice Boltzmann equation. A Chapman-Enskog analysis (or just moment equation Wagner (1997)) of the lattice Boltzmann evolution equation Chapman and Cowling (1990) then recovers the continuity and Navier-Stokes equations of the corresponding lattice Boltzmann model.
For a lattice gas defined by the evolution equation given by Eq. (1), we can find its Boltzmann limit by taking the ensemble average over this evolution equation. We start by defining,
$f_{i}(x,t)=\langle n_{i}(x,t)\rangle,$ (44)
where the non-equilibrium ensemble average delimited-⟨$\langle$..$\rangle$ … ⟩ is defined as,
$\langle X\rangle=\sum_{\{n_{j}\}_{g}}XP(\{n_{j}\}_{g},t),$ (45)
with $P(\{n_{j}\}_{g},t)$ being the probability of finding the system in the specific microstate $\{n_{j}\}_{g}$ at time $t$. We can then take the Boltzmann average of Eq. (1),
$\langle n_{i}(x+v_{i}\Delta t,t+\Delta t)\rangle=\langle n_{i}(x,t)\rangle+\langle\Xi_{i}(\{n_{j}\}_{g})\rangle,$ (46)
which gives,
$f_{i}(x+v_{i}\Delta t,t+\Delta t)=f_{i}(x,t)+\Omega_{i},$ (47)
which is the lattice Boltzmann evolution equation associated with our lattice gas system. $\Omega_{i}$ in this case is our lattice Boltzmann collision operator, which is defined as the Boltzmann average of $\Xi_{i}$,
$\Omega_{i}=\langle\Xi_{i}(\{n_{j}\}_{g})\rangle.$ (48)
We can also derive the $f_{i}$-values for any system for which we know the ensemble average $N$-, $J$-, and $\pi$-values mentioned in the prior section,
$\bar{N}(x,t)=\langle N\rangle,$ (49)
$\bar{N}u=\bar{J}(x,t)=\langle J\rangle,$ (50)
$\bar{\pi}(x,t)=\langle\pi\rangle.$ (51)
This is possible by taking the Boltzmann average in Eq. (45) over Eqs. (9), (10), and (11). These expressions can then be written as,
$f_{1}=\frac{\bar{\pi}+\bar{J}}{2},$ (52)
$f_{0}=\bar{N}-\bar{\pi},$ (53)
$f_{-1}=\frac{\bar{\pi}-\bar{J}}{2}.$ (54)
For our algorithm to have a unique Boltzmann limit, we must be able to derive an expression for $\Omega_{i}$ that depends only on the local $f_{i}$-values. For the Boolean lattice gas this is always possible if the $n_{i}$ are independently distributed, as the $f_{i}$-values then fully define the $n_{i}$ distribution Frisch et al. (1987); Broadwell and Han (2007). When velocity states are allowed to have more than a single particle occupying them, the $f_{i}$-values no longer fully define the $n_{i}$ distribution, even for uncorrelated particle distributions.
For example, we consider the probability distribution defined by a set of delta functions that on average lead to a specific set of $f_{i}$-values,
$P^{\delta}(\{n_{i}\}_{l})=$ (55)
$\prod_{i}(1-f_{i}+\lfloor f_{i}\rfloor)\delta(n_{i}-f_{i})+(f_{i}-\lfloor f_{i}\rfloor)\delta(n_{i}-f_{i}-1),$
which allows us to have a valid delta distributed $\{n_{i}\}_{l}$ state even for non-integer $f_{i}$-values. This ensemble is significantly different to the ensemble recovered from the Poisson distribution in Eq. (31), though their means remain identical.
A unique Boltzmann limit was found for the diffusive sampling lattice gas both by Chopard and Droz (Chopard and Droz, 1998, sec. 5.7) and by Seekins and Wagner Seekins and Wagner (2022). Both groups found that the diffusive sampling lattice gas recovered the BGK lattice Boltzmann collision operator in their Boltzmann limit, given by,
$\Omega_{i}^{\mathrm{BGK}}=-\omega\left(f_{i}-f_{i}^{0}\right).$ (56)
This lattice Boltzmann collision operator sets the particle distribution $\{f_{i}\}_{l}$ to some local equilibrium distribution $\{f_{i}^{0}\}_{l}$, at a rate given by the relaxation time, given by the inverse of our single-timestep particle collision fraction $\omega$. In the Boltzmann limit of the diffusive sampling lattice gas, $f_{i}^{0}$ is the polynomial lattice Boltzmann equilibrium distribution.
Blommel and Wagner showed that the BGK collision operator using the entropic local equilibrium distribution given by Eq. (32) is a good approximation in most cases Blommel and Wagner (2018). In order to test this for our sampling algorithm, we can compare the average collision operators for different distributions by looking at the difference in the average pre- and post-collision $\pi$-values that result from applying the collision operator to those initial distributions.
We also wished to consider the impact of correlations on the collision operator, as highly correlated systems may not have the same Boltzmann limit as uncorrelated systems. Specifically, we considered alternative sets of delta distributed particles with two highly correlated states. In both cases, either all particles at a given lattice site are moving, or all particles at that lattice site are at rest. The probability for a given lattice site to be in a moving state is given by the ratio of $\bar{\pi}$ to $\bar{N}$. For the first probability distribution, the moving particles are evenly distributed within a single lattice site, after taking the initial $\bar{J}$-value into account, giving us,
$P_{\mathrm{corr1}}(\{n_{i}\}_{l};\bar{N},\bar{J},\bar{\pi})=$
$\frac{\bar{\pi}}{\bar{N}}\delta\left(n_{1}-\frac{\bar{N}+\bar{J}}{2}\right)\delta\left(n_{-1}-\frac{\bar{N}-\bar{J}}{2}\right)\delta(n_{0})$
$+\left(1-\frac{\bar{\pi}}{\bar{N}}\right)\delta(n_{1})\delta\left(n_{-1}\right)\delta\left(n_{0}-\bar{N}\right).$ (57)
For the second probability distribution, all lattice sites either have only left moving particles, only right moving particles, or only rest particles, with each lattice site chosen with moving particles having a probability of $0.5+\bar{J}/\bar{N}$ of having only rightward moving particles. Putting this together gives us,
$P_{\mathrm{corr2}}(\{n_{i}\}_{l};\bar{N},\bar{J},\bar{\pi})=$
$\frac{\bar{\pi}}{\bar{N}}\Biggl{(}\left[0.5+\frac{\bar{J}}{\bar{N}}\right]\delta\left[n_{1}-\bar{N}\right]\delta\left[n_{0}\right]\delta\left[n_{-1}\right]+$
$\left[0.5-\frac{\bar{J}}{\bar{N}}\right]\delta\left[n_{-1}-\bar{N}\right]\delta\left[n_{0}\right]\delta\left[n_{1}\right]\Biggr{)}$
$+\left(1-\frac{\bar{\pi}}{\bar{N}}\right)\delta\left(n_{0}-\bar{N}\right)\delta\left(n_{1}\right)\delta\left(n_{-1}\right).$ (58)
As collisions require either a pair of rest particles or a pair of moving particles with opposing velocities, we can expect the number of collisions present in a system with a distribution given by $P_{\mathrm{corr2}}$ to be heavily diminished compared to other distributions, as well as only allowing for an increase in the total $\pi$-value of the system.
To test the Boltzmann limit of our collision operator, we initialized four systems of 100,000 lattice sites utilizing the Poisson distribution given by Eq. (31), the set of delta functions given by Eq. (55), and the two highly correlated distributions $P_{\mathrm{corr1}}$ and $P_{\mathrm{corr2}}$. We started by defining constant (independent of $x$) initial $\bar{N}$- and $\bar{J}$-values and setting the initial $\bar{\pi}$-value to its minimum ($\bar{\pi}=0$). We were then able to define $f_{i}$-values corresponding to those average quantities using Eqs. (52), (53), and (54). Each system was then collided without streaming, collecting each pre- and post-collision set of $\pi$-values, and the average of the resulting difference between the two was taken to be our averaged collision operator.
To compare with the lattice Boltzmann, the BGK collision operator was calculated directly utilizing the $\{f_{i}\}_{l}$ in Eq. (56) and with an equilibrium distribution given by Eq. (32). All simulations were then repeated, increasing the initial $\bar{\pi}$-value by 1 until the $\bar{\pi}$-value reached the $\bar{N}$-value selected.
Figure 2: The averaged collision operator given an imposed average $\pi$-value for four lattice gas particle distributions compared to the BGK collision operator given an initial average ensemble of $\{\bar{N}=30,\bar{J}=0,\bar{\pi}=0\}$. Both the average collision operator (a) and the difference from the BGK calculated by subtracting the BGK from each collision operator (b) are shown. See text for further details.
Fig. 2 shows the results of this test for a series of increasing average $\pi$-values. Fig. 2 (a) shows the average change in $\pi$-value over the course of a single collision, and Fig. 2 (b) shows the difference between each collision operator and the BGK collision operator.
The highly correlated initial condition of Eq. (58) shows a behavior that is notably non-BGK. The equilibrium value of $\pi$ doesn’t correspond to a zero crossing for the collision operator, meaning that $\pi$alues already larger than the equilibrium value will initially grow for this distribution. Clearly this move away from equilibrium will change the distribution so that the $\pi$-value will eventually decrease towards the equilibrium value. However, it is quite unclear how such non-BGK behavior will affect the hydrodynamic transport parameters.
Though the three uncorrelated collision operators and the collision operator resultant from $P_{\mathrm{corr1}}$ are similar, as can be seen in Fig. 2 (a), they are not identical, which Fig. 2 (b) shows more explicitly. We observe an oscillatory behavior of the delta distributed system. This is related to a slightly different average $\pi$-value for even or odd $n_{0}$. The slightly altered slope of the Poisson distributed system from the BGK and delta distributed systems can be interpreted as a slightly altered BGK relaxation value for $\omega$. Importantly here the equilibrium value for $\pi$ is unaffected.
These results show that the Boltzmann limit of our algorithm is not unique. However, for the uncorrelated systems studied, the BGK entropic lattice Boltzmann algorithm is a good approximation. This entropic lattice Boltzmann algorithm is identical to the algorithm first developed by Ansumali et al. Ansumali et al. (2003).
Using the previously discussed Chapman-Enskog analysis, the approximate continuity and Navier-Stokes equations for the D1Q3 sampling lattice gas can be derived as Ansumali et al. (2003),
$\partial_{t}\bar{N}+\partial_{x}\bar{J}=0,$ (59)
and,
$\partial_{t}\bar{J}+\partial_{x}\frac{\bar{J}^{2}}{\bar{N}}=\partial_{x}\bar{N}\theta+\frac{4}{3}\partial_{x}\nu_{\omega}\bar{N}\partial_{x}\frac{\bar{J}}{\bar{N}}.$ (60)
Here $\nu_{\omega}$ is the kinematic viscosity related to the system’s single-timestep particle collision fraction,
$\nu_{\omega}=\theta\left(\frac{1}{\omega}-\frac{1}{2}\right),$ (61)
where $\theta$ is related to the speed of sound on our lattice,
$\theta=c_{s}^{2}=\sum_{i}f_{i}^{eq}(v_{i}-u)^{2}/\sum_{i}f_{i}^{eq}\approx\frac{1}{3}.$ (62)
Blommel and Wagner show in Fig. 5 of their paper that this value for $\theta$ is only a good approximation for $|u|<0.4$ Blommel and Wagner (2018). Outside of that range, $\theta$ will have some dependence on the density, and thus the Navier-Stokes equation will have pressure and viscosity terms that are density dependent.
The kinematic viscosity in Eq. (61) can be lowered by selecting an $\omega$-value between $1$ and $2$, which is useful for simulating low-viscosity systems. This is easily done for the BGK collision operator of Eq. (56). These systems are generally referred to as over-relaxed. However there is no obvious way for the lattice gas to access over-relaxed collisions due to the fact that there is no obvious way to collide more than $100\%$ of particles at a given lattice site. Some work has been done on this problem by Strand and Wagner Strand and Wagner (2022), who created a method to simulate over-relaxation in diffusive systems, but we leave the extension of these methods to hydrodynamic systems to future work.
IV Non-equilibrium Dynamics
IV.1 The Decaying Sound Wave
To test the approximate Boltzmann limit of our lattice gas, we start by analyzing the decaying isothermal sound wave. While the isothermal nature makes this a non-physical application, it is still valuable to compare the lattice Boltzmann and lattice gas algorithms without addressing the complication of including energy conservation. An analytical solution for this system was found by Luiz Czelusniak Czelusniak (2022) in his thesis, and is given by,
$\bar{N}_{\mathrm{sound}}(x,t)$ $=\bar{N}^{eq}+A_{0}\sin\left(\frac{2\tilde{\pi}x}{L}\right)\cos(ft)e^{-\lambda t},$ (63)
$u_{\mathrm{sound}}=$ $B\cos\left(\frac{2\tilde{\pi}x}{L}\right)\sin\left(ft\right)e^{-\lambda t}.$ (64)
where $A_{0}$ is the initial amplitude of the wave. $\lambda$ is the exponential decay rate, which depends on the kinematic viscosity of the system such that,
$\lambda=\left(\frac{2\tilde{\pi}}{L}\right)^{2}\frac{\nu}{\bar{N}^{eq}},$ (65)
$f$ is the oscillation frequency of the sine wave given by,
$f^{2}=\left(\frac{2\tilde{\pi}}{L}\right)^{2}\theta^{2}-\lambda^{2},$ (66)
and $B$ is a constant that depends on the system’s initial conditions,
$B=\frac{-A_{0}\theta}{\bar{N}^{eq}}.$ (67)
Due to the fact that $\pi$ is already defined as a moment in our system, $\tilde{\pi}$ represents the mathematical constant. A key assumption for this analytical solution is that $J$ is small, so that the non-linear term $\partial_{x}\bar{J}^{2}/\bar{N}$ can be dropped in Eq. (60). Thus, we are restricted to small values for $A_{0}$, on the order of $1\%$ of $\bar{N}^{eq}$.
The sound wave must be initialized utilizing a particle distribution for lattice gas systems due to the discrete nature of particles. As the global equilibrium ensemble for the lattice gas is approximately Poisson distributed, we initialize our system by sampling the occupation numbers out of the Poisson distribution from Eq. (31), with $\tilde{f_{i}}$-values given by Eq. (32). The $\bar{N}$- and $u$-values needed to find the $\tilde{f_{i}}$-values can then be defined as the initial conditions for Eqs. (63) and (64) respectively,
$\bar{N}_{\mathrm{sound}}(x,0)$ $=\bar{N}^{eq}+A_{0}\sin\left(\frac{2\tilde{\pi}x}{L}\right),$ (68)
$u_{\mathrm{sound}}(x,0)$ $=0.$ (69)
Figure 3: Normalized pre- and post-collision $\pi$-values of a decaying sound wave. The lattice Boltzmann system was initialized using Eq. (68) with $\bar{N}^{eq}=1,000$ and $L=250$, and the lattice gas used these same values as a mean for its Poisson distributed initialization. After 50 iterations run at $\omega=1$, the pre- and post-collision $\pi$-values were measured over the course of a single collision. The lattice gas was averaged over 100,000 seeds. Both were normalized according to the equilibrium average post-collision $\pi$-value from Eq. (70).
We can observe the effect of a single collision on the system by looking at the difference between the pre- and post-collision averaged $\pi$-values of the system. We then normalize the system with the expected average post-collision $\pi$-value, which is found by using the $f_{i}^{0}$ from Eq. (32),
$\bar{\pi}^{0}=f_{1}^{0}+f_{-1}^{0}.$ (70)
This formula assumes the system is in an equilibrium state, however it serves as a good approximation even for non-equilibrium systems. The results for the effect of a single collision on the decaying sound wave can be seen in Fig. 3, which shows agreement between the average effect of the collision operator for the BGK entropic lattice Boltzmann algorithm from Eq. (56) with an equilibrium distribution given by Eq. (32) and lattice gas algorithm.
To observe the viscosity of the system, we first consider the amplitude decay of the sine wave as it approaches equilibrium. We utilize a method of amplitude extraction discussed by Blommel and Wagner in an appendix to their paper Blommel and Wagner (2018). Specifically, the amplitude of our sound wave is given by,
$A(t)=\frac{\sum_{x}\sin\left(\frac{2\tilde{\pi}x}{L}\right)\bar{N}(x,t)}{\sum_{x}\sin^{2}\left(\frac{2\tilde{\pi}x}{L}\right)}.$ (71)
Figure 4: The extracted amplitude evolution of the decaying sound wave for various values of $\omega$ for the lattice gas and lattice Boltzmann simulations over 850 iterations for a lattice size of 50. Systems were initialized otherwise identically to those in Fig. 3, and run for $\omega=1.0$, $\omega=0.7$, and $\omega=0.1$. The amplitude was extracted utilizing Blommel and Wagner’s method after each iteration Blommel and Wagner (2018). Lattice gas systems were averaged over 25,000 seeds.
Fig. 4 shows the results of this amplitude extraction for $\omega=1$, $\omega=0.7$, and $\omega=0.1$ over 850 iterations for an average of $\bar{N}^{eq}=1,000$ particles. There is excellent agreement between the lattice gas and the lattice Boltzmann algorithms in this case, despite the fact that the BGK lattice Boltzmann algorithm is not an exact Boltzmann limit for our lattice gas. We can measure the viscosity of the system numerically by extracting the decay rate from the wave amplitude. We can do this by fitting our results from Fig. 4 to the analytical solution from Eq. (63). The amplitude of Eq. (63) is given by,
$A^{th}(t)=A_{0}\cos(ft)e^{-\lambda t}.$ (72)
The recovered $\lambda$ can then be converted to the kinematic viscosity $\nu$ by inverting Eq. (65).
Figure 5: The recovered viscosities of the lattice Boltzmann and lattice Gas corresponding to different values of $\omega$ along with the theoretical scaling from Eq. (61). Each system was initialized and run identically to those in Fig. 4. Along with the systems from Fig. 4, systems were also run at $\omega=0.3$ and $\omega=0.5$. The $\lambda$-values were found via non-linear curve fitting on xmgrace with the analytical solution from Eq. (72) and calculated using Eq. (65).
We can then compare the extracted values of $\nu$ from the lattice Boltzmann and lattice gas systems to the theoretical $\nu_{\omega}$-value from Eq. (61) for a given value of $\omega$. Fig. 5 shows the results of this comparison for a high density ($\bar{N}^{eq}=1,000$) system.
Figure 6: The extracted amplitude of the decaying sound wave for both lattice gas and lattice Boltzmann simulations over 850 iterations run at $\omega=1$. along with two potential fittings at a constant decay rate of for the lattice gas. Simulations are initialized identically to those in Fig. 4, except with a $\bar{N}^{eq}$-value of 10. Both fits have been made utilizing non-linear curve fitting on xmgrace assuming Eq. (72) as the fitting equation, with the fits restricted to Iterations$<300$ (early-time) and $400<$Iterations$<700$ (late-time).
There is excellent agreement between the results of the lattice Boltzmann and lattice gas algorithms in this case, as well as between the results from both algorithms and Eq. (61). However, this is only true when considering high density systems. For low density systems, the decay rate is non-constant, an effect that can be seen in Fig. 6.
Both the lattice gas and its fittings show a clear distinction from the lattice Boltzmann in the case of this smaller $\bar{N}^{eq}$-value. The fits themselves are also imperfect, due to the fact that Eq. (72) assumes $\lambda$ to be a constant value. However because the early-time fitting and late-time fitting do not agree with each other, the value of $\lambda$ is not constant with respect to time for the lattice gas.
Figure 7: The density dependence of the viscosity of a decaying sound wave for lattice Boltzmann and lattice gas systems. See text for details.
By extracting the decay rate and calculating the viscosity using Eq. (65), as was done in Fig. 5, and repeating the process for various values of $\bar{N}^{eq}$, we can visualize the scaling of the lattice gas’s viscosity with respect to density. This is shown in Fig 7.
We start by initializing the systems identically to those of Fig. 4, but with varying initial values of $\bar{N}^{eq}$. All systems were then run with $\omega=1$ for 1,000 timesteps, averaged over 100,000 seeds per density. The viscosities were calculated identically to those in Fig 5. Two sets of fitting parameters were used, corresponding to the early-time and late-time fittings in Fig. 6. We then repeated this process for the BGK entropic lattice Boltzmann algorithm.
We also considered the collision operator for an uncorrelated Poisson distribution. Since the slope of the collision operator from Fig. 2 (a) is slightly different to that of the BGK collision operator, we can obtain a different effective $\omega$-value with which to calculate the viscosity. However, as we will see below, this small correction does not explain the much larger deviations measured in our simulation.
There is a clear deviation from the theoretical viscosity for lower densities in both the early-time and late-time fittings for the lattice gas. The lattice gas’s late-time fittings in particular show increased deviation from the theoretical viscosity value even for higher densities. The lattice Boltzmann agrees well with the theoretical viscosity value for both early-time and late-time fittings. This reinforces our findings from Fig. 6 that there is a difference in the non-equilibrium behavior of the lattice Boltzmann and lattice gas algorithms that leads to this non-constant decay rate. However as the density increases beyond approximately $\bar{N}^{eq}=1,000$, the difference becomes small, particularly for early-time fittings.
Both lattice gas fittings having a more extreme viscosity scaling compared to that of the collision operator alone implies that the system is not perfectly uncorrelated as it relaxes to equilibrium. An example of this effect was already shown in Fig. 2 (a). The inconsistency in the decay rates at low densities indicates that there is some underlying physical effect that leads to a buildup of correlations that decrease the efficiency of the collision operator leading to a smaller effective collision probability. Studying this effect in detail, however, is outside of the scope of this paper, and will be left to future work.
IV.2 The SOD Shock Tube
The SOD shock tube is defined by an initial state with two regions in equilibrium with differing initial average densities, but identical initial velocities Sod (1978). Due to the step discontinuity between the two regions, the SOD shock tube is a Riemann problem Toro (2009). This particular Riemann problem has been used as a basic test case for a variety of hydrodynamic simulation methods Sod (1978); Zha (1999); Loubère et al. (2014); Feng et al. (2010).
In our case, we define the initial distribution of average densities as,
$\bar{N}_{\mathrm{shock}}(x,0)=\begin{cases}N_{l}&x<\frac{L}{2}\\ N_{r}&\frac{L}{2}\leq x\\ \end{cases}$ (73)
with $N_{l}$ and $N_{r}$ being the average densities of the two initial regions. The initial $\tilde{f}_{i}$-values for the shock tube can then be written using Eq. (32), as both regions should initially be in equilibrium internally. The boundary conditions typically include reflective boundary conditions on both ends of the tube, however we found that a mirrored shock wave system allowed us to use periodic boundary conditions, though only one side of the shock wave will be shown due to their symmetry.
The analytical solution for the SOD shock tube for zero viscosity is relatively well studied as a solution to the Euler equations Toro (2009), however our collision operator restricts the system to an isothermal average state. Negro et al. derived an analytical solution for the isothermal inviscid SOD shock tube system Negro et al. (2019). For a rightwards moving shock ($N_{l}>N_{r}$) Negro’s analytical solutions for density and velocity can be written as,
$\bar{N}_{\mathrm{shock}}(x,t)=\begin{cases}N_{l}&x<x_{r}\\ N_{l}\exp\left(\frac{x_{0}-x}{tc_{s}}+1\right)&x_{r}\leq x<x_{c}\\ N_{l}\exp(-\zeta)&x_{c}\leq x<x_{s}\\ N_{r}&x\geq x_{s},\\ \end{cases}$ (74)
and,
$u_{\mathrm{shock}}(x,t)=\begin{cases}0&x<x_{r}\\ \frac{x-x_{0}}{t}+c_{s}&x_{r}\leq x<x_{c}\\ \zeta c_{s}&x_{c}\leq x<x_{s}\\ 0&x\geq x_{s},\end{cases}$ (75)
with the assumption that $u(x,0)=0$ for all $x$. $x_{s}$ is the position of the shock wave front, and is given by,
$x_{s}=\frac{\zeta c_{s}t}{1-\frac{N_{r}}{N_{l}}\exp(\zeta)}+x_{0}.\$ (76)
We can also find $x_{c}$, the start of the central density plateau, and $x_{r}$, the start of the rarefaction fan,
$x_{c}$ $=c_{s}(\zeta-1)t+x_{0},$ (77)
$x_{r}$ $=x_{0}-c_{s}t.$ (78)
$x_{0}$ is the initial position of the discontinuity between regions, and for our initial conditions is given by $x_{0}=L/2$. $\zeta$ is a dimensionless quantity that can be found by solving,
$02+\zeta^{2}-\frac{N_{l}}{N_{r}}\exp(-\zeta)-\frac{N_{r}}{N_{l}}\exp(\zeta)=$ 0 . (79)
In a finite system, this analytical solution only remains valid until the end of the boundary is reached by the wave in a system with reflective boundary conditions, or until the two shock wave fronts meet for our mirrored system with periodic boundary conditions.
Figure 8: The $\bar{N}$ (a) and $\bar{J}$ (b) moment graphs of the isothermal SOD shock tube at 75 iterations for $\omega=1$ comparing the lattice gas results averaged over 25,000 seeds and the lattice Boltzmann results. The system was initialized with a size of 150 lattice sites, and $\bar{N}=\rho$ values of $N_{l}=1,000$ and $N_{r}=500$, along with a mirrored shock wave on the opposite side of the periodic boundary. The lattice gas $n_{i}$-values are assigned utilizing a Poisson distribution around Eq. (32), with the $\bar{N}_{\mathrm{shock}}$-values specified previously and $u_{\mathrm{shock}}(x,0)=0$. (a) includes the analytical solution from Eq. (74). The analytical solution in (b) is the product of Eqs. (74) and (75). (a) also includes labels for various features of the shock wave.
Fig. 8 shows the equivalence between the lattice Boltzmann and lattice gas simulations with the analytical solution after evolution. Specifically, it considers the ensemble averaged values of the $N$ and $J$ moments: $\bar{N}$ and $\bar{J}$ from Eqs. (49) and (50). $\bar{N}$ can be compared directly with $\bar{N}_{\mathrm{shock}}$ from the analytical solution in Eq. (74), and $\bar{J}$ can be compared to the analytical solutions in Eqs. (74) and (75) multiplied together, since $\bar{J}=u\bar{N}$. Fig. 8 (a) also labels the features of the SOD shock wave, including the rarefaction fan, the shock wave front, and the positions $x_{r}$, $x_{c}$, and $x_{s}$.
The averaged $N$ (Fig. 8 (a)) and $J$ (Fig. 8 (b)) moments for the lattice gas algorithms show an excellent agreement with those of the lattice Boltzmann algorithm. They also show decent agreement with their respective analytical solutions for the inviscid SOD shock tube. There are some small deviations between both sets of simulated data and the analytical solution due to the viscosity not being low enough to recover the piecewise nature of the analytical solutions. For lattice Boltzmann simulations, it is possible to achieve closer agreement by using $\omega>1$, however, as mentioned above, collision probabilities greater than one are not yet an option for the hydrodynamic lattice gas algorithm.
While gathering data for Fig. 8, we also considered the initial fluctuations in the moments, as we wished to see whether initial system fluctuations had any effect on the collision operator the way correlations were shown to in Fig. 2 (a). To this end, simulations with two different initial particle distributions were run. The first of these distributions is a Poisson distribution of the $n_{i}$-values given by Eq. (31), with Eq. (73) and $N_{l}=1,000,N_{r}=500$ for its initial $\bar{N}$-values. The second distribution was a flat distribution given by a delta function around the entropic lattice Boltzmann equilibrium distribution,
$P_{SOD\delta}(\{n_{i}\}_{l},x,0)=\prod_{i}\delta\left(n_{i}-f_{i}^{0}\left(\bar{N}(x,0),u(x,0)\right)\right).$ (80)
$P_{SOD\delta}(\{n_{i}\}_{l},0)$ also has the restriction that $f_{i}^{\mathrm{0}}\in\mathbb{N}_{0}$. As the averaged values of the moments resultant from both of these initial probability distributions are indistinguishable to the order of $10^{-4}$, which is within the bounds of the averaging of our data set, only the results from the Poisson distributed system are shown for the lattice gas in Fig. 8.
Figure 9: Normalized pre- and post-collision $\pi$-values of the SOD Shock Tube. The systems were initialized identically to those in Fig. 8. After 75 iterations, another collision was run for each system and the pre- and post-collision $\pi$-values were measured for that extra collision. The lattice gas was averaged over 100,000 seeds. Both were normalized according to the average post-collision $\pi$-value from Eq. (70).
Since the system dynamics are driven by the collision operator we examined the pre- and post-collision $\bar{\pi}$-values for a single collision, similarly to what was done in Fig. 3. We can see in Fig. 9 that these pre- and post-collision results show good agreement between the lattice Boltzmann and lattice gas algorithms. We can also note the fact that for the parts of this system that are in equilibrium, the average collision operator is identical between the pre- and post-collision states. These equilibrium regions include both those that were initialized in equilibrium and have yet to be affected by the shock behavior, as well as the bulk of the shock wave itself ($x_{c}\leq x<x_{s}$). The other regions of the wave, including the rarefaction fan and the shock front, show highly non-equilibrium behavior. The shock front shows an increase in the average pre-collision $\pi$-value, and the rarefaction fan shows a decrease in the average pre-collision $\pi$-value.
IV.3 Deviation from the Analytical Solution
Looking at how the lattice Boltzmann and the lattice gas algorithms behave under extreme initial conditions for the SOD shock tube system allows us to compare how both algorithms deviate from the analytical solution. As the analytical solution found by Negro is a solution to the Euler equations, it is necessarily inviscid, and thus assumes that the viscosity of the system is zero Negro et al. (2019). Therefore, the analytical solution for the SOD shock tube should differ from the data from both algorithms to some degree for any finite value of $\omega<2$, as can be seen in Fig. 8.
Figure 10: The $\bar{N}$ (a), $\bar{J}$ (b), and $\bar{\pi}$ (c) moment graphs of the isothermal SOD shock tube at 75 iterations for $\omega=0.1$ comparing the lattice gas results averaged over 25,000 seeds and the lattice Boltzmann results. All systems were initialized and run identically to those in Fig. 8 aside from the differing $\omega$-value. (a) and (b) include identical analytical solutions to those in Fig 8. (c) includes the equilibrium $\bar{\pi}^{0}$ given by Eq. (70).
Fig. 10 compares the lattice Boltzmann and lattice gas results for a system with a much higher viscosity, being run at $\omega=0.1$. Specifically it compares the $\bar{N}$ (Fig. 10 (a)), $\bar{J}$ (Fig. 10 (b)), and post-collision $\bar{\pi}$ (Fig. 10 (c)) moments to the applicable analytical solutions in Fig. 10 (a) and (b), and the equilibrium average $\bar{\pi}$-value in Fig. 10 (c).
Though the agreement with the analytical solutions for $\bar{N}$ and $\bar{J}$ is poor due to the large difference in viscosity between the simulations and the analytical solution, the lattice gas and lattice Boltzmann results show excellent agreement with each other. Similarly, the fact that $\omega=0.1$ means that the system will not fully relax to equilibrium in one timestep, meaning that the post-collision $\bar{\pi}$-values from the lattice Boltzmann and lattice gas algorithms do not recover the equilibrium $\bar{\pi}$-value, however the two simulations agree well with each other. This is important to note, as it implies that even the highly non-equilibrium state that the SOD shock tube represents still has a Boltzmann limit that is very well approximated by the BGK entropic lattice Boltzmann algorithm, something that Fig. 2 shows is not guaranteed.
To further test the deviation from the analytical solution of the lattice Boltzmann and lattice gas results, we can consider an extreme initial density disparity between $N_{l}$ and $N_{r}$. This leads to a deviation from the analytical solution because as the initial density disparity between the two regions increases, the analytical solution requires $u_{\mathrm{shock}}$-values that approach or exceed $1$. Because $1$ is the maximum possible value of $u$ a lattice Boltzmann or lattice gas system can have, simulations in this range typically show significant deviations. To show this, we consider the extreme density ratio of $20:1$ for $N_{l}$ and $N_{r}$. For this particular density ratio, the $u_{\mathrm{shock}}$-value of the analytical solution is approximately $0.9$ on the central density plateau, as opposed to the analytical solution for the density ratio from Fig. 8, which has a $u_{\mathrm{shock}}$-value on the central density plateau of only around $0.2$.
Figure 11: The $\bar{N}$ (a) and $\bar{J}$ (b) moment graphs of the isothermal SOD shock tube at 75 iterations for $\omega=1$ comparing the lattice gas results averaged over 25,000 seeds and the lattice Boltzmann results. The systems were initialized with $\bar{N}$ values of $N_{l}=1,000$ and $N_{r}=50$, but were otherwise initialized and run identically to those in Fig. 8. Analytical solutions calculated identically to those in Fig. 8, aside from the different $N_{l}$ and $N_{r}$ values used, are included for $\bar{N}$ and $\bar{J}$.
Fig. 11 shows the results of this increased density disparity for the $\bar{N}$ (Fig. 11 (a)) and $\bar{J}$ (Fig. 11 (b)) moments, as well as including the applicable analytical solutions. The agreement between the lattice Boltzmann and lattice gas systems is excellent, as it was in both Fig. 8 and Fig. 10, however as expected the agreement between the simulation results and the analytical solution is poor.
Looking at the velocity on the plateau of the wave ($x_{c}\leq x<x_{s}$) which is where it is maximal, there is a clear difference between the simulation results and the analytical solution. The $\bar{J}$-value on the plateau corresponds to a $u_{\mathrm{shock}}$-value of approximately $0.78$ for the lattice Boltzmann and lattice gas simulations. This velocity disparity also explains why the values of $x_{c}$ and $x_{s}$ differ between the simulated data and the analytical solutions, as the $\theta$-values for the high velocity state will not be $1/3$, leading to artifacts in the Navier-Stokes equation as was discussed previously Blommel and Wagner (2018).
IV.4 Non-equilibrium Fluctuations in the Isothermal SOD Shock Tube
While fluctuations in equilibrium are reasonably well understood, much less is known about non-equilibrium fluctuations. Our integer lattice gas algorithm allows us to examine these non-equilibrium fluctuations. Thus, we consider here the non-equilibrium fluctuations of an isothermal SOD shock wave as a proof of principle.
We limit our investigation of the non-equilibrium fluctuations to the squared standard deviation of a quantity $X$,
$\sigma_{X}^{2}=\left\langle(X-\langle X\rangle)^{2}\right\rangle.$ (81)
Utilizing the same set of simulations from Fig. 8, we examined the fluctuation of the moments in the lattice gas, comparing their average non-equilibrium fluctuations to those of global equilibrium. As the equilibrium fluctuations in $N$- and $\pi$-values are given by Eq. (31), we normalize the average non-equilibrium fluctuations for $N$ and $\pi$ with $\bar{N}$ and $\bar{\pi}$ respectively.
The equilibrium fluctuations in the $J$-moment do not have the local $\bar{J}$-value as their average, as $J$ is represented by the difference between two Poisson distributed quantities (those being $n_{1}$ and $n_{-1}$). Therefore, its global equilibrium fluctuations are Skellam distributed Skellam (1946). A Skellam distribution has a variance corresponding to the sum of the mean values of the two Poisson distributions that make up the Skellum distribution Jose (1956). The average value for our $J$-fluctuations, therefore, becomes $\bar{\pi}$.
Figure 12: Normalized nonequilibrium fluctuations of the $N$, $J$, and $\pi$ moments after 75 iterations. Initialized identically to the system in Fig. 8, the system was run for 75 iterations at $\omega=1$ and averaged over 25,000 seeds. Each moment’s variance was then calculated and normalized by dividing by its mean. For the $N$-moment this mean is the local $\bar{N}$-value (Fig. 8(a)) and for the $J$ and $\pi$ moments this mean is the $\bar{\pi}$-value from Eq. (70).
The fluctuation results can be seen in Fig. 12. In the bulk of the shock wave and the regions initialized in equilibrium, the non-equilibrium fluctuations match the equilibrium fluctuations, deviating slightly due to limited averaging. However, in the highly non-equilibrium parts of the shock wave the non-equilibrium fluctuations clearly deviate from the equilibrium fluctuations, with a slight decrease of the fluctuations in the rarefaction fan and a drastic increase around the shock front, mirroring the effects seen in the pre-collision $\pi$-values in Fig. 9. While this correlation is very evocative we don’t yet understand why it occurs.
The slight decrease in the average $J$ fluctuations in the bulk of the shock wave when compared to the equilibrium fluctuations was noticed as a potential aberration. However when tested in a flat system with an initial advected velocity it was determined to be a product of the $J$ fluctuations taking additional time to reach their equilibrium value rather than a directional bias in the algorithm.
Figure 13: Normalized nonequilibrium $N$-moment fluctuations for Poisson and delta distributed initial fluctuations after 75 iterations. The Poisson distributed system was initialized identically to the system in Fig. 8, however the flat system was initialized utilizing constant $n_{i}$ values corresponding to the values of $f^{0}_{i}$ as in Eq. (80). The systems were run, and the variances were calculated and normalized identically to those in Fig. 12.
This delayed timescale for the buildup of fluctuations is also present when the system is initialized with a flat density distribution corresponding to Eq. (80). Since there are no initial fluctuations, all moment fluctuations must build up over time, leading to reduced fluctuations compared to those present in Fig. 12. These reduced fluctuations can be seen in Fig 13.
IV.5 Galilean Invariance of the SOD Shock Tube
As stated previously, the non-Galilean invariance of the collision operator for Boolean lattice gases was a major factor in the widespread adoption of the BGK lattice Boltzmann model Frisch et al. (1987); Qian et al. (1992). Due to this, it is important that we show to what extent our lattice gas algorithm maintains Galilean invariance if we seek the sampling lattice gas algorithm to be competitive with the lattice Boltzmann algorithm.
In order to numerically test the Galilean invariance of our system, we initialized a set of SOD shock tube systems similarly to those in Fig. 8 for both the lattice Boltzmann and lattice gas algorithms with a lattice size of $L=1,000$ and a net velocity of zero, and a second set of identical systems with a non-zero constant velocity offset $u_{g}$. After running all simulations (including averaging 2,500 seeds for each lattice gas system) for 400 iterations at $\omega=1$, we adjusted the coordinate systems for the initial $u_{g}$-values to take into account the additional distance the simulation would have traveled on average due to the velocity by using,
$x^{\prime}=[x+\lfloor u_{g}t\rfloor]\bmod[{L}],$ (82)
which takes into account our system’s periodic boundary conditions. $[a]\bmod[b]$ is the modulo operator, returning the positive remainder when $b$ divides $a$. We then calculated the mean squared error in $\bar{N}$ and $u$,
$\epsilon(\bar{N}(t))=\frac{1}{L}\sum^{L}_{x=0}(\bar{N}(x^{\prime},t)-\bar{N}_{0}(x,t))^{2},$ (83)
$\epsilon(u(t))=\frac{1}{L}\sum^{L}_{x=0}(u(x^{\prime},t)-u_{g}-u_{0}(x,t))^{2},$ (84)
where $\bar{N_{0}}$ and $u_{0}$ are the $\bar{N}$ and $u$ data for the $u_{g}=0$ system.
Because we used a mirrored shock wave that propagates opposite the shock waves shown in the previous figures in place of a wall, we can look at the effect of both positive and negative initial values of $u_{g}$. To do this, the lattice was divided at the center of the shock wave, and the shock moving in the positive x-direction was used to define the error for $u_{g}>0$, and the shock moving in the negative x-direction was used to define the error for $u_{g}<0$. Thus, the error for the shock moving in the positive x-direction is higher than the error in the negative x-direction because it is moving at a faster velocity relative to $u_{g}$.
Figure 14: Mean Square Error for both the lattice gas and lattice Boltzmann for $\bar{N}$ (a) and $u$ (b). See text for details.
A comparison of the $\epsilon$-values for $\bar{N}$ and $u$ for both the lattice gas and lattice Boltzmann algorithms are shown in Fig. 14 (a) and 14 (b) for various values of $u_{g}$ ranging from $-0.2$ to $0.2$. The mean squared error for the lattice gas algorithm is on an order of magnitude comparable to that of the lattice Boltzmann algorithm for both $\bar{N}$ and $u$, meaning that the lattice gas and lattice Boltzmann algorithms are approximately equivalent in terms of Galilean invariance. There is a slight deviation observed for higher values of $u_{g}$, which implies that the lattice gas algorithm is slightly more Galilean invariant than the lattice Boltzmann algorithm, however we are as of yet unsure why this deviation occurs.
V Computational Efficiency
As the main drawback of Blommel and Wagner’s method was its lack of computational efficiency, determining the efficiency of our own method is necessary for comparison. This first requires some more technical discussion of the methods we used to make our sampling method more efficient. Because the local equilibrium ensemble out of which we wish to sample is not a well known distribution, using a library such as the GNU Scientific Library (GSL) in C is non-trivial Galassi et al. (2018). Though packages such as the GSL have the ability to sample from arbitrary distributions, the full distribution must be known in advance, which is not the case for our system. This is due to the fact that the $N$- and $J$-values that define the local $\pi$ equilibrium ensemble are not known until after the streaming step of the prior timestep. The ideal solution to this issue would be a library of full distributions pre-calculated for every combination of $N$- and $J$-values, however this method requires either a large amount of memory (scaling with $O(N)$) or to read the distribution from a file for each collision.
Thus, we decided to utilize a dynamic lookup table. This method is similar to what was done in our prior paper on the diffusive sampling lattice gas, only with the added need to manage the $J$-value as well as the $N$-value Seekins and Wagner (2022). We first initialized the lookup table as a two-dimensional triangular array with $12N-$ 1 entries in each sub-array, with a likely maximum amount of particles that would be at one lattice site given by three times the average system density. Should a state occur beyond this maximum density, the table would be rebuilt with a higher maximum density in the algorithm. As the system was running, the approximate center point of the cumulative equilibrium distribution ($\sum_{\pi=J}P^{0}(\pi;N,J)\approx 0.5$) of the local equilibrium ensemble was saved in the lookup table. This then allowed us to then navigate to any part of the local equilibrium ensemble by utilizing the recursive formula from Eq. (27) and its inverse. Because this dynamic lookup table adds entries on the fly, over the course of longer runs generating new table entries will become rare. A copy of the code for this algorithm is available on Github at Seekins (2025).
Even with this lookup table, the limiting factor of efficiency for the lattice gas is the generation of sampled random numbers. In order to investigate the maximum possible improvement in efficiency due to this sampling, we wished to determine the speed of the algorithm without the sampling as a factor. To do this, we pre-generated a full set of random numbers for a simulation, and then fed those into the system in place of the dynamic lookup table.
Refer to caption (b)
Figure 15: Computational efficiency of the lattice Boltzmann algorithm compared to various lattice gas algorithms for average system densities from $N=10$ to $N=1,000$. (a) compares the lattice Boltzmann algorithm to Blommel and Wagner’s algorithm, our sampling lattice gas algorithm, and the maximal potential efficiency of our sampling lattice gas algorithm for $\omega=0.5$. (b) compares the lattice Boltzmann algorithm to our sampling lattice gas algorithm and its maximal efficiency for $\omega=1$, leaving out Blommel and Wagner’s algorithm since it has no true way to reach $\omega=1$. See text for further details.
We ran our efficiency tests for four separate algorithms: a lattice Boltzmann algorithm using the entropic BGK collision operator, Blommel and Wagner’s algorithm, our sampling lattice gas algorithm, and the maximal lattice gas algorithm. The results of these comparisons are shown in Fig. 15 in millions of lattice updates per second (MLUPS). Because Blommel and Wagner’s algorithm is based around individual two-particle collisions, the particle collision fraction can be estimated by taking Seekins and Wagner (2022),
$C=N\ln(1-\omega).$ (85)
This equation can only approximate a system with $\omega\neq 1$, as to approach $\omega=1$ would require an infinite number of collisions. Thus, we only compare our algorithm to Blommel and Wagner’s algorithm for $\omega\neq 1$.
In the case of Fig. 15 (a), we compare Blommel’s algorithm to the lattice Boltzmann, the sampling lattice gas, along with the maximum possible efficiency of the sampling lattice gas for various average densities, all initialized with a lattice size of $L=100$ and run for $\omega=0.5$. The starting average density is initialized via Eq. (32), for flat $\bar{N}(x,0)$ and $u(x,0)=0$. The lattice gas systems are initialized using a Poisson distribution around these $f_{i}^{eq}$ values given by Eq. (31). The sampling lattice gas systems are run for $t=10,000$ timesteps to allow the computational cost of initializing the lookup tables to become negligible.
All systems are then timed for 10,000 timesteps, and are transformed into MLUPS by taking,
$t_{MLUPS}=\frac{Lt}{1,000,000t_{s}},$ (86)
where $t_{s}$ is the time taken to run in seconds. The simulations were run on a Dell XPS laptop with an Intel I5 CPU, and a clock speed of 1.6 GHz.
For the maximal lattice gas, a full sampling lattice gas system was run, saving the sampled $\pi$-values and binomial random numbers (for $\omega\neq 1$) into arrays, before repeating the simulation and reading all random variables from those arrays. Because Blommel and Wagner’s algorithm takes far longer to run than the others, we ran a linear regression on a smaller sample of data to extrapolate the efficiency of the algorithm. The explicitly measured data points are displayed, along with the linear regression itself, and due to the linear scaling, a logarithmic scale is necessary on the y-axis for proper resolution.
The sampling lattice gas and lattice Boltzmann timings are on the same order of magnitude, though the lattice gas is approximately a factor of five slower. The maximum efficiency for the lattice gas algorithm, however, is nearly a factor of three faster than the efficiency of the lattice Boltzmann algorithm. This indicates that the sampling lattice gas algorithm has the potential to surpass the lattice Boltzmann algorithm in efficiency should the sampling algorithm for the $\pi$-distribution and binomial distribution be improved enough. All of the aforementioned algorithms are much more efficient than Blommel and Wagner’s algorithm.
Fig 15 (b) repeats the above test for all algorithms aside from the Blommel and Wagner algorithm for $\omega=1$. Due to the lack of additional binomial sampling needed for an $\omega=1$ system, the sampling lattice gas algorithm is more efficient, and thus only ends up approximately a factor of two worse than the lattice Boltzmann algorithm, and the maximal efficiency of the lattice gas is just over a factor of three better. Due to Fig. 15 (b) not needing to account for Blommel and Wagner’s algorithm, a linear scale is able to be used for the y-axis. Both graphs show that our sampling lattice gas algorithm is indeed competitive with the lattice Boltzmann algorithm, and may become superior in terms of direct efficiency should it be further optimized in the future.
VI Outlook
Integer lattice gases are a promising simulation method for fluctuating systems, and they are also directly derivable from Molecular Dynamics simulations via the MDLG method Parsa and Wagner (2017). This is why we were motivated to look into them more closely. They also form a physical microsystem that has a Boltzmann limit that can lead to novel insights into developing unconditionally stable lattice Boltzmann methods. Lattice gas methods generally contain additional physical content like correlations, that cannot be captured with current lattice Boltzmann methods, as we saw in section III. This is why we intend to continue to develop integer lattice gas approaches.
There are several remaining open problems when considering the development of the sampling lattice gas algorithm. The most immediately practical of these is the extension to two and three dimensional systems. This would allow the full recovery of the capabilities of the Blommel and Wagner collision operator, and would also allow deeper investigations into the scaling effects of the spuriously conserved parity of the $n_{0}$-values. The main hurdle for this development is in finding the proper method for distributing momentum across multiple axes while maintaining conservation.
Over-relaxation is a further open question, and is one with some progress towards its development. As discussed in Section III, Strand and Wagner have developed a method for simulating over-relaxation in the diffusive case Strand and Wagner (2022). However, this method involves a momentum-flipping operator, and thus applying it to hydrodynamic systems is non-trivial.
Adding forcing to the collision operator is another direction of potential study, which would allow the simulation of non-ideal systems. This addition may utilize the coarse-graining of a molecular dynamics simulation, as was done by Parsa and Wagner previously for the ideal lattice gas, to analyze the form of the forcing lattice gas collision operator Parsa and Wagner (2017).
In terms of studying the algorithm itself, the non-uniqueness of the collision operator and the nature of the method’s fluctuations in equilibrium and non-equilibrium systems warrant additional attention. Along with these two, the issue of potential non-equilibrium correlations and their effect on relaxation at low densities is another open problem to be considered.
Finally, further optimization of the sampling lattice gas algorithm is needed in order to truly match the efficiency of or outperform the lattice Boltzmann algorithm for hydrodynamic simulations. Though the lattice Boltzmann algorithm may have the advantage of removing the noise from simulations inherently, those systems that require natural fluctuations would find lattice gas algorithms to be a superior algorithm for practical simulation.
Acknowledgements.
We would like to thank Matteo Lulli for discussing with us the potential influence of correlations on the collision operator, as well as ways to analyze this correlated state more thoroughly for future work. We would also like to thank Victor Ambrus for pointing out the isothermal SOD shock tube solution Negro et al. (2019).